OpenStudy (anonymous):

plz help. I need an answer quick. give ya medal. Find the maximum value of the function f(x)= -x^2+3x+2. a. 4.250 b.1.500 c.there is no maximum ( was my choice but wrong) d.cannot be determined

OpenStudy (campbell_st):

find the vertex of the parabola... and easy way is to find the line of symmetry \[x = \frac{-b}{2a}\] then substitute that value into the equation... to get the max value. the reason you use the line of symmetry is that the max(or min) is on the line of symmetry in your question a = -1 and b = 3

OpenStudy (anonymous):

so my answer will be choice B? or A?

OpenStudy (campbell_st):

so if you calculate the line of symmetry, you get x = 3/2 so substitute into the equation to find the max value for y

OpenStudy (anonymous):

I still do not get the answer. it is not in any of my choices. I already tried choices b and c. both are wrong. and I am running out of time.

OpenStudy (campbell_st):

well this is the calculation \[y = -(\frac{3}{2})^2 + 3 \times \frac{3}{2} + 2\]

OpenStudy (anonymous):

so it would be choice a? this is my last chance to get this question right.

OpenStudy (campbell_st):

well if that's what the calculation says then that's correct

OpenStudy (anonymous):

I thought you knew which answer it was and you were guiding me to it.